OpenStudy (anonymous):

Grade 8 math please help! will give medal for answer.. what is: -5x + 4 = 34. x is?

OpenStudy (tkhunny):

Have you considered subtracting 4 from each side?

OpenStudy (anonymous):

hmmm, no I haven't 1 second let me try it!

OpenStudy (anonymous):

yeah, I'm not sure..heh heh..

OpenStudy (tkhunny):

What do you get when you subtract 4 from 4? What do you get when you subtract 4 from 34?

OpenStudy (anonymous):

0 and 30

OpenStudy (tkhunny):

Okay, this transforms the equation... -5x + 4 = 34 Becomes -5x + 4 - 4 = 34 - 4 -5x + 0 = 30 or just -5x = 30 What do you suppose we should do next?

OpenStudy (anonymous):

ummmm do something with numbers, haha! idk

OpenStudy (tkhunny):

Not good enough. You must get 'x' all by itself. What is it that is now the problem with that goal? Is 'x' all alone?

OpenStudy (anonymous):

do we multiply the 5 and 30? or add?

OpenStudy (tkhunny):

How is the "-5" attached to the 'x'. What operation is implied by writing "-5" and 'x' next to each other like that?

OpenStudy (anonymous):

uhhh, okay sorry I don't know, that's why I'm on here. I suck at math!

OpenStudy (tkhunny):

Hogwash. If I write "3 dogs" What does that mean about the number of dogs I have?

OpenStudy (here_to_help15):

you will divide.

OpenStudy (tkhunny):

Grrrr.....

OpenStudy (here_to_help15):

;)

OpenStudy (here_to_help15):

Keep in mind that if you divide a positive and a negative you will get a negative, but if you divide a negative and a negative you get a positive.

OpenStudy (here_to_help15):

Number

OpenStudy (kidrah69):

@Oak what is -5x/-5=? 30/-5=?

OpenStudy (anonymous):

thanks guys!!!

OpenStudy (kidrah69):

Answer the question :P

OpenStudy (here_to_help15):

Np :) Mostly a great thanks to @tkhunny of course :)

OpenStudy (anonymous):

6

OpenStudy (here_to_help15):

Incorrect

OpenStudy (kidrah69):

No remeber its -5

OpenStudy (here_to_help15):

Remember the rule i told you above

OpenStudy (kidrah69):

negative / positive = negative

OpenStudy (here_to_help15):

@kidrah69 :) next time *Talking very gently* let him answer :D

OpenStudy (anonymous):

ohhhh right sorry!

OpenStudy (kidrah69):

@Here_to_Help15 i made it stand out :P

OpenStudy (kidrah69):

So @Oak what will be our answer?

OpenStudy (here_to_help15):

lol i personally would have waited and reminded him about the rule with the - and + 's,expecting him to say the answer

OpenStudy (anonymous):

-6...?

OpenStudy (kidrah69):

correct! :D

OpenStudy (kidrah69):

oh and also :) \(\huge\color{orange}{WELCOME~TO~OpenStudy!!~:)}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (here_to_help15):

OOOOOps @kidrah69 i am a idiot my apologies i thought you told him the answer ;)

OpenStudy (anonymous):

awesome thankyou guys!!! you are all s awesome!!!

OpenStudy (kidrah69):

:-)

OpenStudy (here_to_help15):

*walks away slowly*